# DualDynamics: Synergizing Implicit and Explicit Methods for Robust Irregular Time Series Analysis

### v3

@@ Introduction
Effectively modeling time series data is a cornerstone in machine learning, supporting numerous applications across diverse sectors. This significance has driven substantial research endeavors aimed at developing novel methodologies capable of accurately capturing the complexities of continuous latent processes (Chen et al. 2018; Kidger et al. 2020; Oh, Lim, and Kim 2024).
Recent advancements in this field have led to two distinct approaches: implicit methods, such as Neural Differential Equation (NDE)-based techniques, and explicit methods, exemplified by Neural Flows. Neural Ordinary Differential Equations (Neural ODEs) (Chen et al. 2018), Neural Controlled Differential Equations (Neural CDEs) (Kidger et al. 2020) and Neural Stochastic Differential Equations (Neural SDEs) (Oh, Lim, and Kim 2024) have emerged as prominent implicit methods, recognized for their ability to learn continuous-time dynamics and underlying temporal structures. These approaches directly learn continuous latent representations based on vector fields parameterized by neural networks. However, NDE-based methods face challenges including limited expressive power (Dupont, Doucet, and Teh 2019; Massaroli et al. 2020b; Chalvidal et al. 2021; Kidger 2022) and scalability issues when analyzing irregular or complex time series data (Norcliffe et al. 2020; Morrill et al. 2021, 2022; Irie, Faccio, and Schmidhuber 2022; Pal et al. 2023). These limitations are attributed to data complexity, sequence length variations, and the stability constraints of numerical solvers (He and Semnani 2023; Westny et al. 2023).
In parallel, explicit methods such as Neural Flows have been developed, focusing on directly mapping solution curves using neural networks (Lu et al. 2018; Sonoda and Murata 2019; Massaroli et al. 2020a; Biloš et al. 2021). These approaches offer invertible solutions and enhanced stability through the change of variables formula (Kobyzev, Prince, and Brubaker 2020; Papamakarios et al. 2021). However, they struggle with irregularly-sampled time series and are sensitive to initial state and corresponding initial value problem.
@@
The landscape of neural differential equations for time series modeling has been dominated by two distinct approaches: implicit methods, exemplified by Neural ODEs, Neural CDEs and Neural SDEs, and explicit methods, such as Neural Flows. These approaches fundamentally differ in how they represent and solve the underlying dynamics of time-evolving systems.
@@
Let ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$ be a vector of original irregularly-sampled observations and ${\bm{z}}(t)$ denote a latent state at time $t$. Continuous latent dynamics ${\bm{z}}(t)$ can be achived by integration of ${\textnormal{d}}{\bm{z}}(t)$ over time in $[0,t]$. Neural Differential Equations can be expressed as:
@@
with initial condition ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ is an affine function. $f$ and $g$ are neural networks parameterized by $\theta_{f}$ and $\theta_{g}$, respectively. $X(t)$ is a controlled path for Neural CDEs, and $W(t)$ is a Wiener process for Neural SDEs.
@@
Model architecture
To illustrate our DualDynamics framework, we utilize Neural CDEs as the primary example in this section. Neural CDEs serve as an effective representation of the implicit component in our dual approach, offering a robust foundation for modeling irregular time series data. Figure 1 shows the conceptual overview of the proposed architecture using Neural CDE (implicit) and Flow-based model (explicit) for classification task.
@@
where $f^{*}(s;\theta_{f},X):=f(s,{\bm{z}}(s);\theta_{f})\frac{{\textnormal{d}}X(s)}{{\textnormal{d}}s}$. Recall that $X$ is generated from ${\bm{x}}$. Here, we implement $X$ for ‘Hermite Cubic splines with a backward difference’, which is smooth and unique (Morrill et al. 2022). Through this part, we find ${\bm{z}}\in\mathbb{R}^{d_{z}}$, and then plug it into the following flow model ${\mathcal{G}}$ as:
@@
It is crucial for ensuring that the total probability across transformations remains constant, enabling accurate modeling of dynamic systems’ evolution without loss or gain of information (Li and Chen 2008; Dinh, Sohl-Dickstein, and Bengio 2016). For each fixed $t$, $p_{{\bm{z}}(t)}$ and $p_{\hat{{\bm{z}}}(t)}$ be the probability density functions of ${\bm{z}}(t)$ and $\hat{{\bm{z}}}(t)$, respectively. From the instantaneous change of variables formula (Chen et al. 2018), we observe that
@@
where the last inequality is obtained by changing the order of partial derivatives. Incorporating the flow model ${\mathcal{G}}$ into the original NDE-based model enhances its expressive power without significantly increasing the computational burden, as opposed to the addition of an extra ODE/CDE formulation.
@@
To optimize parameters involved in the proposed method, we employ adjoint-based backpropagation (Chen et al. 2018; Kidger et al. 2020; Kidger 2022). This method allows for efficient gradient computation, ensuring that the information from one step influences and refines the subsequent step, leading to a more unified optimization process. The adjoint method is a powerful technique often used to calculate the gradients of systems governed by differential equations. The basic idea is to recast the derivative computation problem as the adjoint equation (Pontryagin 2018; Pollini, Lavan, and Amir 2018).
@@
To estimate the model parameters $\theta_{f}$ and $\theta_{\mathcal{G}}$, we compute the gradients of the loss function with respect to these parameters:
@@
All experiments were performed using a server on Ubuntu 22.04 LTS, equipped with an Intel(R) Xeon(R) Gold 6242 CPU and a cluster of NVIDIA A100 40GB GPUs. The source code for our experiments can be accessed at https://github.com/yongkyung-oh/DualDynamics.
@@
The implicit component in our framework primarily utilizes Neural CDE, with Neural ODE and Neural SDE explored as alternatives in our ablation study. For the explicit component, we evaluated three distinct flow models: ResNet Flow, GRU Flow, and Coupling Flow. These flow architectures, as outlined by Biloš et al. (2021), were incorporated into our proposed framework. To assess the impact of using flow models, we also implemented a conventional multilayer perceptron (MLP) as a baseline for comparison. In practice, optimal flow model is selected by hyperparameter tuning process.
@@
We employed a variety of benchmark models for classification and forecasting, incorporating conventional RNN (Rumelhart, Hinton, and Williams 1986; Medsker and Jain 1999), LSTM (Hochreiter and Schmidhuber 1997), and GRU (Chung et al. 2014), alongside their notable variants, including GRU-$\Delta t$ (Choi et al. 2016), and GRU-D (Che et al. 2018). Also, we used attention-based method including transformer (Vaswani et al. 2017), Multi-Time Attention Networks (MTAN) (Shukla and Marlin 2021), and Multi-Integration Attention Module (MIAM) (Lee et al. 2022). ODE-based models like, GRU-ODE (De Brouwer et al. 2019), ODE-RNN (Rubanova, Chen, and Duvenaud 2019) ODE-LSTM (Lechner and Hasani 2020), Latent-ODE (Rubanova, Chen, and Duvenaud 2019), Augmented-ODE (Dupont, Doucet, and Teh 2019), and Attentive co-evolving neural ordinary differential equations (ACE-NODE) (Jhin et al. 2021) are considerd. Furthermore, Neural CDE (Kidger et al. 2020), and Neural Rough Differential Equation (Neural RDE) (Morrill et al. 2021), were included in our comparative study. Additionally, recent advances in Neural CDE are considered, including Attentive Neural Controlled Differential Equation (ANCDE) (Jhin et al. 2024), EXtrapolation and InTerpolation-based model (EXIT) (Jhin et al. 2022), and LEArnable Path-based model (LEAP) (Jhin et al. 2023). Lastly, variations of Neural Flow (Biloš et al. 2021) are also included.
@@
Table 1: Average classification performance on 18 datasets under regular and three missing rates (Values in parentheses show the average of 18 standard deviations. The top-ranking and runner-up choices are emphasized.
@@
To provide comprehensive insights, we conducted a descriptive analysis of model performances using the ‘BasicMotions’ dataset as a case study. We focused on five key methods: a standard RNN, GRU-ODE, ODE-LSTM, Neural CDE, Neural Flow, and our proposed method. Our investigation covered four different settings, including regular (0% missingness), and three missing rates including 30%, 50%, and 70%.
(a) Classification performance comparisons
(b) Test loss without missingness (regular)
(c) Test loss with 50% missing (irregular)
Figure 2: Performance and stability comparison using ‘BasicMotions’ dataset (monitoring for 100 epochs)
Figure 2 (a) illustrates the performance trend of each model concerning varying missing data rates. While the conventional RNN’s performance decreases as the missing rate increases, GRU-ODE and ODE-LSTM outperform conventional RNNs but become less stability with increasing missing data. Neural CDE demonstrates superior performance, whereas Neural Flow shows performance worse than that of the conventional RNN.
Our proposed method consistently demonstrates remarkable performance regardless of missing data rates. Further exploration is presented in Figure 2 (b) and Figure 2 (c), illustrating the test loss dynamics of the ’BasicMotions’ dataset, both with and without missingness. Despite its simplicity, Neural Flow shows limited convergence in both cases, whereas our proposed method achieves stability, outperforming both conventional approaches and Neural CDE. Furthermore, we included detailed results and computational time analysis in Supplementary material.
Table 2: Ablation study of the proposed method (Average of 18 standard deviations in parentheses)
@@
Based on our initial findings, we extended our experiments to 18 public benchmark datasets from the the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository (Bagnall et al. 2018; Dau et al. 2019), encompassing Motion & Human Activity Recognition (HAR), Electrocardiogram (ECG) & Electroencephalogram (EEG), and Sensor domains. Following Kidger et al. (2020) and Oh, Lim, and Kim (2024), we generated subsets with 30%, 50%, and 70% missingness rates, creating four distinct scenarios. Data was split 70:15:15 for training, validation, and testing. We employed five-fold cross-validation, evaluating mean performance and ranking across folds. For dataset details and implementation specifics, refer to the Supplementary material and the corresponding repository [^3] https://github.com/yongkyung-oh/torch-ists.
@@
We conducted an in-depth comparative analysis of the classification efficacy and cross-entropy loss across various flow model designs, as detailed in Table 2. For implicit component, Neural CDE outperformed compared to Neural ODE and Neural SDE. The proposed method selected the flow configuration from ResNet flow, GRU flow, and Coupling flow during hyperparameter tuning. For the comparison, we employed conventional Multi-layer Perceptrons (MLPs), which notably failed to meet specific flow criteria, including invertibility. Empirical evaluations indicate that the performance of conventional MLP is worse compared to the flow-based approach. This observation confirms the superiority of our proposed method.
Interpolation for the missing data
The interpolation experiment utilized the 2012 PhysioNet Mortality dataset (Silva et al. 2012), comprising 37 variables from ICU records over 48 hours post-admission. Following Rubanova, Chen, and Duvenaud (2019), timestamps were adjusted to the nearest minute, yielding up to 2880 intervals per series. We adhered to Shukla and Marlin (2021)’s methodology, varying observation frequency from 50% to 90% to predict remaining samples using 8000 instances. In our interpolation experiments, we followed the experimental protocol advocated by Shukla and Marlin (2021). This strategy involved generating interpolated values based on a selected subset of data points to predict values for the unobserved time intervals. The interpolation spanned an observational range from 50% to 90% of the total data points. During the test phase, models were designed to interpolate values for the remaining intervals in the test set. The efficacy of the models was quantitatively assessed using Mean Squared Error (MSE), supported by five cross-validations.
Table 3 demonstrates that the interpolation performance of our proposed methods is superior compared to the benchmark models across all three flow configurations. The flow configuration achieved lower MSE errors compared to the non-flow neural network, a consistent and robust observation across various levels of observed data, ranging from 50% to 90%. Detailed experiment protocols and results are included in Supplementary material.
@@
{adjustwidth}
-1em-1em Methods Test AUROC Memory (MB) OI No OI OI No OI GRU-$\Delta t$ 0.878 ± 0.006 0.840 ± 0.007 837 826 GRU-D 0.871 ± 0.022 0.850 ± 0.013 889 878 GRU-ODE 0.852 ± 0.010 0.771 ± 0.024 454 273 ODE-RNN 0.874 ± 0.016 0.833 ± 0.020 696 686 Latent-ODE 0.787 ± 0.011 0.495 ± 0.002 133 126 ACE-NODE 0.804 ± 0.010 0.514 ± 0.003 194 218 Neural CDE 0.880 ± 0.006 0.776 ± 0.009 244 122 ANCDE 0.900 ± 0.002 0.823 ± 0.003 285 129 EXIT 0.913 ± 0.002 0.836 ± 0.003 257 127 Proposed method 0.918 ± 0.003 0.873 ± 0.004 453 233
Table 4: AUROC on PhysioNet Sepsis dataset
@@
Table 5: Forecasting performance on MuJoCo dataset
@@
Figure 3: Forecasting performance versus the number of parameters on MuJoCo with regular scenario
Table 6: Forecasting performance on Google dataset
@@
Figure 4: Forecasting performance versus the number of parameters on Google with regular scenario

### v4

@@ Abstract
Codehttps://github.com/yongkyung-oh/DualDynamics
Extended versionhttps://arxiv.org/abs/2401.04979
@@ Introduction
Effectively modeling time series data is a cornerstone in machine learning, supporting numerous applications across diverse sectors. This significance has driven substantial research endeavors aimed at developing novel methodologies capable of accurately capturing the complexities of continuous latent processes (Chen et al. 2018; Kidger et al. 2020; Kidger 2022; Oh, Lim, and Kim 2024).
Recent advancements have led to two distinct approaches: implicit methods, such as Neural Differential Equation (NDE)-based techniques, and explicit methods, such as Neural Flows. Neural Ordinary Differential Equations (Neural ODEs) (Chen et al. 2018), Neural Controlled Differential Equations (Neural CDEs) (Kidger et al. 2020) and Neural Stochastic Differential Equations (Neural SDEs) (Oh, Lim, and Kim 2024) have emerged as prominent implicit methods, recognized for their ability to learn continuous-time dynamics and underlying temporal structures. These approaches directly learn continuous latent representations based on vector fields parameterized by neural networks. However, NDE-based methods face challenges including limited expressive power (Dupont, Doucet, and Teh 2019; Massaroli et al. 2020b; Chalvidal et al. 2021; Kidger 2022) and scalability issues when analyzing irregular or complex time series data (Norcliffe et al. 2020; Morrill et al. 2021, 2022; Irie, Faccio, and Schmidhuber 2022; Pal et al. 2023). These limitations are attributed to data complexity, sequence length variations, and the stability constraints of numerical solvers (He and Semnani 2023; Westny et al. 2023).
In parallel, explicit methods have been developed, focusing on directly mapping solution curves using neural networks (Lu et al. 2018; Sonoda and Murata 2019; Massaroli et al. 2020a; Biloš et al. 2021). These approaches offer invertible solutions and enhanced stability through the change of variables formula (Kobyzev, Prince, and Brubaker 2020; Papamakarios et al. 2021). However, they struggle with irregularly-sampled time series and are sensitive to initial state and corresponding initial value problem.
@@
The landscape of neural differential equations for time series modeling has been dominated by two distinct approaches: implicit methods, exemplified by Neural ODEs, Neural CDEs and Neural SDEs, and explicit methods, such as Neural Flows. These approaches fundamentally differ in how they represent and solve the underlying temporal dynamics.
@@
Let ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$ be a vector of original irregularly-sampled observations and ${\bm{z}}(t)$ denote a latent state at time $t$. Continuous latent dynamics ${\bm{z}}(t)$ can be achieved by integration of ${\textnormal{d}}{\bm{z}}(t)$ over time in $[0,t]$. Neural Differential Equations can be expressed as:
@@
with initial condition ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ is an affine function. $f$ and $g$ are neural networks parameterized by $\theta_{f}$ and $\theta_{g}$, respectively. $X(t)$ is a control path for Neural CDEs, and $W(t)$ is a Wiener process (or Brownian motion) for Neural SDEs.
@@
Proposed Framework
To illustrate our DualDynamics framework, we utilize Neural CDEs as the primary example in this section. Neural CDEs serve as an effective representation of the implicit component in our dual approach, offering a robust foundation for modeling irregular time series data.
Figure 1 shows the conceptual overview of the proposed architecture using Neural CDE (implicit) and Flow-based model (explicit) for classification task.
@@
where $f^{*}(s;\theta_{f},X):=f(s,{\bm{z}}(s);\theta_{f})\frac{{\textnormal{d}}X(s)}{{\textnormal{d}}s}$. Recall that $X$ is generated from ${\bm{x}}$. Here, we implement $X$ for ‘Hermite Cubic splines with a backward difference’, which is smooth and unique, as discussed by Morrill et al. (2022). Through this part, we find ${\bm{z}}\in\mathbb{R}^{d_{z}}$, and then plug it into the following flow model ${\mathcal{G}}$ as:
@@
It is crucial for ensuring that the total probability across transformations remains constant, enabling accurate modeling of dynamic systems’ evolution without loss or gain of information (Li and Chen 2008; Dinh, Sohl-Dickstein, and Bengio 2016). For each fixed $t$, $p_{{\bm{z}}(t)}$ and $p_{\hat{{\bm{z}}}(t)}$ be the probability density functions of ${\bm{z}}(t)$ and $\hat{{\bm{z}}}(t)$, respectively. From the instantaneous change of variables formula (Chen et al. 2018), we observe that:
@@
where the last inequality is obtained by changing the order of partial derivatives. Incorporating the flow model ${\mathcal{G}}$ into the original NDE-based model enhances its expressive power without significantly increasing the computational burden, as opposed to the addition of an extra NDE formulation.
@@
To optimize parameters involved in the proposed method, we employ adjoint-based backpropagation (Chen et al. 2018; Kidger et al. 2020; Kidger 2022; Oh, Lim, and Kim 2024). This method allows for efficient gradient computation, ensuring that the information from one step influences and refines the subsequent step, leading to a more unified optimization process. The adjoint method is a powerful technique often used to calculate the gradients of systems governed by differential equations. The basic idea is to recast the derivative computation problem as the adjoint equation (Pontryagin 2018; Pollini, Lavan, and Amir 2018).
@@
To estimate the model parameters $\theta_{f}$ and $\theta_{\mathcal{G}}$, we compute the gradients of the loss function with respect to these parameters simultaneously:
@@
All experiments were performed using a server on Ubuntu 22.04 LTS, equipped with an Intel(R) Xeon(R) Gold 6242 CPU and a cluster of NVIDIA A100 40GB GPUs. The source code for our experiments and datasets can be accessed at https://github.com/yongkyung-oh/DualDynamics. Please refer the supplementary material for the experiment settings and implementation details.
@@
The implicit component in our framework primarily utilizes Neural CDE, with Neural ODE and Neural SDE explored as alternatives in our ablation study. For the explicit component, we evaluated three distinct flow models: ResNet Flow, GRU Flow, and Coupling Flow. These flow architectures, as outlined by Biloš et al. (2021), were incorporated into our proposed framework. To assess the impact of using flow models, we also implemented a conventional multilayer perceptron (MLP) as a baseline for comparison. In practice, optimal flow model is selected by hyperparameter tuning stage.
@@
In experiments, we employed a variety of benchmark models for classification and forecasting, incorporating conventional RNN (Rumelhart, Hinton, and Williams 1986; Medsker and Jain 1999), LSTM (Hochreiter and Schmidhuber 1997), and GRU (Chung et al. 2014), alongside their notable variants, including GRU-$\Delta t$ (Choi et al. 2016), and GRU-D (Che et al. 2018). Also, we used attention-based method including transformer (Vaswani et al. 2017), Multi-Time Attention Networks (MTAN) (Shukla and Marlin 2021), and Multi-Integration Attention Module (MIAM) (Lee et al. 2022). ODE-based models like, GRU-ODE (De Brouwer et al. 2019), ODE-RNN (Rubanova, Chen, and Duvenaud 2019) ODE-LSTM (Lechner and Hasani 2020), Latent-ODE (Rubanova, Chen, and Duvenaud 2019), Augmented-ODE (Dupont, Doucet, and Teh 2019), and Attentive co-evolving neural ordinary differential equations (ACE-NODE) (Jhin et al. 2021) are considerd. Furthermore, Neural CDE (Kidger et al. 2020), and Neural Rough Differential Equation (Neural RDE) (Morrill et al. 2021), were included in our comparative study. Additionally, recent advances in Neural CDE are considered, including Attentive Neural Controlled Differential Equation (ANCDE) (Jhin et al. 2024), EXtrapolation and InTerpolation-based model (EXIT) (Jhin et al. 2022), and LEArnable Path-based model (LEAP) (Jhin et al. 2023). Lastly, variations of Neural Flow (Biloš et al. 2021) are also included.
@@
Table 1: Average classification performance on 18 datasets under regular and three missing rates (Values in parentheses show the average of 18 standard deviations. The top-ranking and runner-up choices are emphasized.)
@@
Descriptive Example.
To provide comprehensive insights, we conducted a descriptive analysis of model performances using the ‘BasicMotions’ dataset as a case study. We focused on five key methods: a standard RNN, GRU-ODE, ODE-LSTM, Neural CDE, Neural Flow, and our proposed method. Our investigation covered four different settings, including regular (0% missingness), and three missing rates including 30%, 50%, and 70%, respectively.
Figure 2: Classification performance comparison on the ‘BasicMotions’ dataset with different scenarios
Figure 2 illustrates the classification performance trend of each model concerning varying missing data rates. While the conventional RNN’s performance decreases as the missing rate increases, GRU-ODE and ODE-LSTM outperform conventional RNNs but become less stability with increasing missing data. Neural CDE demonstrates superior performance, whereas Neural Flow shows performance worse than that of the conventional RNN.
(a) Test loss without missingness (regular)
(b) Test loss with 50% missing (irregular)
Figure 3: Stability comparison on the ‘BasicMotions‘ dataset, monitored over 100 epochs without early stopping
Our proposed method consistently demonstrates remarkable performance regardless of missing data rates. Further exploration is presented in Figure 3 (a) and (b), illustrating the test loss dynamics of the ‘BasicMotions’ dataset, both with and without missingness. Despite its simplicity, Neural Flow shows limited convergence in both cases, whereas our proposed method achieves stability, outperforming both conventional approaches and Neural CDE. Furthermore, we included detailed results and computational time analysis in the supplementary material.
@@
Comprehensive Analysis.
Based on our initial findings, we extended our experiments to 18 public benchmark datasets from the the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository (Bagnall et al. 2018; Dau et al. 2019), encompassing Motion & Human Activity Recognition (HAR), Electrocardiogram (ECG) & Electroencephalogram (EEG), and Sensor domains. Following Kidger et al. (2020) and Oh, Lim, and Kim (2024), we generated subsets with 30%, 50%, and 70% missingness rates, creating four distinct scenarios. Data was split 70:15:15 for training, validation, and testing. We employed five repetition, evaluating mean performance and ranking across iterations.
@@
We conducted an in-depth comparative analysis of the classification efficacy and cross-entropy loss across various flow model designs, as detailed in Table 2. For implicit component, Neural CDE outperformed compared to Neural ODE and Neural SDE. The proposed method selected the flow configuration from ResNet flow, GRU flow, and Coupling flow during hyperparameter tuning. For the comparison, we employed conventional MLP, which notably failed to meet specific flow criteria. Empirical evaluations indicate that the performance of conventional MLP is worse compared to the flow-based approach. This observation confirms the superiority of our proposed method.
Interpolation of Missing Data
The interpolation experiment utilized the 2012 PhysioNet Mortality dataset (Silva et al. 2012), comprising 37 variables from ICU records over 48 hours post-admission. Following Rubanova, Chen, and Duvenaud (2019), timestamps were adjusted to the nearest minute, yielding up to 2880 intervals per series. We adhered to Shukla and Marlin (2021)’s methodology, varying observation frequency from 50% to 90% to predict remaining samples using 8000 instances.
In our interpolation experiments, we followed the experimental protocol advocated by Shukla and Marlin (2021). This strategy involved generating interpolated values based on a selected subset of data points to predict values for the unobserved time intervals. The interpolation spanned an observational range from 50% to 90% of the total data points. During the test phase, models were designed to interpolate values for the remaining intervals in the test set. The efficacy of the models was quantitatively assessed using Mean Squared Error (MSE), supported by five cross-validations.
Table 3 demonstrates that the interpolation performance of our proposed methods is superior compared to the benchmark models across all three flow configurations. The flow configuration achieved lower MSE errors compared to the non-flow neural network, a consistent and robust observation across various levels of observed data, ranging from 50% to 90%. Detailed experiment protocols and additional results are included in the supplementary material.
@@
Methods Test AUROC Memory (MB)
OI No OI OI No OI
GRU-$\Delta t$ 0.878 ± 0.006 0.840 ± 0.007 837 826
GRU-D 0.871 ± 0.022 0.850 ± 0.013 889 878
GRU-ODE 0.852 ± 0.010 0.771 ± 0.024 454 273
ODE-RNN 0.874 ± 0.016 0.833 ± 0.020 696 686
Latent-ODE 0.787 ± 0.011 0.495 ± 0.002 133 126
ACE-NODE 0.804 ± 0.010 0.514 ± 0.003 194 218
Neural CDE 0.880 ± 0.006 0.776 ± 0.009 244 122
ANCDE 0.900 ± 0.002 0.823 ± 0.003 285 129
EXIT 0.913 ± 0.002 0.836 ± 0.003 257 127
Proposed method 0.918 ± 0.003 0.873 ± 0.004 453 233
Table 4: AUROC on PhysioNet Sepsis dataset
@@
We utilized two datasets, MuJoCo and Google, with different levels of observation for the forecasting task. Detailed explanations of each dataset, implementation details, and results are provided in the supplementary material.
MuJoCo.
We utilized the MuJoCo dataset (Todorov, Erez, and Tassa 2012), based on the Hopper configuration from the DeepMind control suite (Tassa et al. 2018). Following Jhin et al. (2023, 2024), we used 50 initial points to predict the next 10, introducing 30%, 50%, and 70% missing data scenarios. Table 5 illustrates the forecasting performance across diverse conditions, with our method consistently demonstrating minimal MSE in all settings.
@@
Table 5: Forecasting performance comparison on the MuJoCo dataset with varying percentages of data dropped
Figure 4: Forecasting performance versus the number of parameters on MuJoCo dataset with regular scenario
Google.
@@
Table 6: Forecasting performance comparison on the Google dataset with varying percentages of data dropped
Figure 5: Forecasting performance versus the number of parameters on Google dataset with regular scenario
Discussion
Performance versus Network Capacity.
Our analysis reveals that the performance improvements of DualDynamics are not merely a result of increased network capacity, but rather stem from the synergistic integration of implicit and explicit methods in our framework.
To validate DualDynamics’ effectiveness, we conducted controlled experiments comparing it against models with similar parameter counts. Figures 4 and 5 illustrate the relationship between parameter count and MSE for our flow architecture versus conventional MLPs. Figure 4 shows that increasing MLP capacity yields limited improvement on the MuJoCo dataset, while DualDynamics maintains superior performance. Figure 5 demonstrates that with optimal flow configuration, DualDynamics consistently outperforms MLPs on the Google dataset across all model sizes.
This performance advantage persists even when adjusting baseline model complexity to match DualDynamics, indicating that our framework’s superiority stems from its architectural design rather than network capacity. These results highlight DualDynamics’ efficiency in capturing complex temporal dynamics in irregular time series.